# MalCVE: Malware Detection and CVE Association Using Large Language Models

### v1

@@ 5.3.4. Reliability
The experiments were restricted by limited resources, totaling approximately \$58.75 for OpenAI services and \$69.86 for Gemini across 3,839 analyzed JARs. Despite the low costs, the limited budget constrained the scale and depth of the experiments relative to typical commercial research standards. Larger and more diverse experiments, which are often feasible in industry-funded studies, were infeasible within the scope of our study.
@@ 6. Conclusion and future work
Malware can cause significant societal consequences. This study investigated the use of decompilation and deobfuscation paired with LLMs to detect whether binary files are malicious. The approach involved decompiling and deobfuscating the code, after which an LLM summarized the behavior and provided a verdict. The solution achieved a 97% mean accuracy across 3,839 analyzed JAR binaries, offering a low-cost, low-effort alternative to dynamic analysis tools, fine-tuning approaches, and commercial platforms. We also investigated the use of LLMs and RAG to associate malware with the CVEs it may potentially exploit. The results demonstrate that LLMs supported by deobfuscation and semantic similarity search can satisfactorily associate malware with CVEs, although there remains room for improvement in accuracy.
@@ 7. Data availability
The evaluation results (in JSON files) are available at:
http://bit.ly/45einJm

### v2

@@ 5.3.4. Reliability
The experiments were restricted by limited resources, totaling approximately \$58.75 for OpenAI services and \$69.86 for Gemini across 3,839 analyzed JARs. We did not compare the accuracy of MalCVE with commercial tools due to their high cost. Our limited budget constrained the scale and depth of the experiments relative to typical commercial research standards. Larger and more diverse experiments, which are often feasible in industry-funded studies, were infeasible within the budget and scope of our study.
@@ 6. Conclusion and future work
Malware can cause significant societal consequences. This study investigated the use of decompilation and deobfuscation paired with LLMs to detect whether binary files are malicious. The approach involved decompiling and deobfuscating the code, after which a LLM summarized the program’s behavior and rendered a verdict. The solution achieved a 97% mean accuracy across 3,839 analyzed JAR binaries, offering a low-cost, low-effort alternative to dynamic analysis tools, fine-tuning approaches, and commercial platforms. We also investigated the use of LLMs and RAG to associate malware with the CVEs it may potentially exploit. The results demonstrate that LLMs supported by deobfuscation and semantic similarity search can satisfactorily associate malware with CVEs, although there remains room for improvement in accuracy.
@@ 7. Data availability
The software and data replication packages are published at:
https://doi.org/10.5281/zenodo.16233415